Solve the following right triangle by finding all of the missing sides and angles. Round your answers to the nearest hundredth. B = 83.7°, C = 90°, and b = 3.21 km
Answer by Boreal(15235)

tan B=b/a
9.0579=3.21/a
a=0.354 km
A=6.3 degrees
C is the sqrt of both of those sides squared and added= sqrt (10.43)=3.23 km